### v1

# PDFMathTranslate: Scientific Document Translation Preserving Layouts

Abstract
Language barriers in scientific documents hinder the diffusion and development of science and technologies. However, prior efforts in translating such documents largely overlooked the information in layouts. To bridge the gap, we introduce PDFMathTranslate, the world’s first open-source software for translating scientific documents while preserving layouts. Leveraging the most recent advances in large language models and precise layout detection, we contribute to the community with key improvements in precision, flexibility, and efficiency. The work has been open-sourced at https://github.com/byaidu/pdfmathtranslate with more than 22k downloads.
PDFMathTranslate: Scientific Document Translation Preserving Layouts
[…]
1 Introduction
The advances in science and technologies spread and evolve in different languages  Von Gizycki ; Montgomery . Yet, language barriers hinder the diffusion and development of scientific progress Ramírez-Castañeda (2020); Hwang (2005); Ammon (2012). For example, while 98% publications in science are written in English Liu ; Ammon , the language only has 7.3% native speakers and no more than 20% speakers worldwide Bahji et al. . Consequently, the majority of human beings are hindered by language barriers from the advances in science and technology. To overcome the barrier, there are attempts such as those from international organizations UNESCO and academic efforts in improving Machine Translation (MT) over text Brown et al. (a); Vaswani et al. (2017); Zhu et al. ; Johnson et al. ; Sennrich et al. .
However, machine translation literature fails to address the unique challenges posed by the layouts in technical translation Schubert (2012). In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures have rich and important meanings. Those text-based attempts ignore the information in the layout and are thus insufficient to address the barrier in scientific and technical document translation.
Figure 1: The architecture of PDFMathTranslate
To fill the gap, we introduce PDFMathTranslate, the world’s first open-source tool to translate PDF documents with preserved layouts. By leveraging recent advances in layout detection and large language models (see Figure 1), we better address the barrier with at least five key contributions: (1) efficient workflow of layout detection, translation, and re-rendering; (2) support for multiple languages (3) support for multiple translation models and services; (4) diverse user interfaces; and (5) a community-commerce model affording sustainable developments.
2 Architecture and Design
Overall, PDFMathTranslate is designed to translate documents while preserving their original layouts. First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service (see Appendix A). Next, it detects the layouts to extract the layouts and textual contents from the document. Third, the texts are translated using the selected translation service, such as GPT-4, DeepL, Google, or Ollama (see Appendix A). Finally, the translated text and previously detected layouts are re-rendered as a translated document with preserved layouts.
Technically, we design the architecture for precision, flexibility, and efficiency. Specifically, we introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, to be explained in the following three sub-sections.
Precise layout parser
Preserving layouts prerequires precise parsing. To precisely parse the layouts, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering. In the beginning, we exploit a recent advance in layout detection,  DocLayout-YOLO-DocStructBench-onnx. The model is another version of a SOTA solution in object detection, Yolov10 Wang et al. , and is fast and accurate in the specific task Zhao et al. .
To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX) onnx standard and a pytorch version. The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardwares.
Flexible translation middleware
Translation middleware offers flexibility in terms of language diversity, supported services, and customization capabilities. One, PDFMathTranslate supports at least 56 languages (see Appendix A), including popular ones such as English, and some languages shared by relatively smaller communities, such as Cherokee. Users can freely select any of these languages as either the input or output language when translating documents.
Two, PDFMathTranslate supports at least 23 popular services for translation. The list of services includes prominent online services such as Google Translate and OpenAI, as well as local deployments of models for either translation or dialogs (see Appendix A). In addition to explicitly listed services, users can integrate any new service complying with OpenAI protocols  [^1] See: https://platform.openai.com/docs/api-reference/chat/create.
Moreover, to simplify the addition of new services, we designed a structure that separates layout flows from translation flows. In this structure, language services receive only textual inputs, enhancing the generalizability and sustainability of our project. As a result, implementing any language service requires only a concise function of fewer than 15 lines.
Three, PDFMathTranslate provides additional customizability in translation. We designed a customization feature prompting strategies that potentially increase the quality of translation or its adaptability in domain-specific tasks. This feature allows users to exploit advanced prompting strategies such as few-shots Brown et al. (b), Chain-of-Thought Wei et al. , Role-playing Shanahan et al. , etc.
Efficient streaming flow
Translating scientific and technical documents, in some cases, needs an efficient workflow for thousands of pages, challenging the performance. To address the issue, we landed on a streaming design that is centered around an efficient in-memory processing pipeline that minimizes disk I/O while preserving document fidelity. Specifically, the translate_stream function accepts a byte stream of a PDF file and transforms it into a mutable document representation. This approach allows the system to dynamically apply modifications, such as converting to PDF/A format when needed and embedding the appropriate fonts based on the target language. By employing temporary buffers and leveraging functions like download_remote_fonts, the design effectively manages resources and ensures that the original file remains unaltered during processing.
Our workflow supports various translation backends, with translators configured as plugins within our workflow. Different translators integrate distinct translation functionalities by inheriting from a base class and implementing the do_translate method. These translators can be seamlessly integrated into different user interfaces. With contributions from the community, our workflow currently supports more than 20 translators. Additionally, it provides robust extensibility for backend LLM services with the OpenAI-compatible API and traditional machine translation service.
Further optimization in performance is achieved through a combination of asynchronous execution, caching strategies, and error-handling mechanisms. Asynchronous constructs, such as asyncio.Event, are employed to support task cancellation and concurrency, particularly when processing large documents or handling multiple translation tasks simultaneously. In addition, the design incorporates a caching mechanism—controlled via the ignore_cache flag—to prevent redundant computations, save LLM tokens, and accelerate the translation process. While a subset font embedding strategy minimizes the final file size without compromising compatibility. Overall, this streaming design reflects a robust, scalable, and resource-efficient framework tailored for automated PDF translation workflows.
3 Usage and Deployments
PDFMathTranslate has been implemented in various interfaces, including a command-line tool (CLI), a graphic user interface (GUI), cross-platform applications (on MacOS and Windows), and Docker images. In addition to those officially supported interfaces, user guaguastandup contributes a Zotero plugin to our community.
For laymen, we offer GUI on Mac  [^2] See https://github.com/reycn/pdf2zh-mac., Windows, Web, online demos, and Zotero plugin  [^3] See https://github.com/guaguastandup/zotero-pdf2zh; community-contributed.. Specifically, desktop versions on Mac and Windows can be installed, demos are publicly accessible online, and GUI can be started using pdf2zh -i. Those graphic interfaces are shown in Figure 2.
Figure 2: Officially supported interfaces
Figure 3: Major commands for CLI
For developers, we support CLI and Docker for advanced usage and substream developments. Developers who prefer command-line tools can conveniently install our Python CLI program using pip install pdf2zh. The command-line tool supports extensive features, with the major ones shown in Figure 3 and documented in detail  [^4] See https://github.com/Byaidu/PDFMathTranslate/blob
/main/docs/ADVANCED.md
. For users aiming to deploy applications on servers, we offer a Docker image designed specifically to simplify online production deployment. Those images are built and distributed on different platforms for developers with potential network issues.
docker pull byaidu/pdf2zh
docker run -d -p 7860:7860 byaidu/pdf2zh
OURS IMT5 Doc2X TeX based6 Google DeepL
Accessibility open-source
Deployment
API integration
Price Free Paid Paid Free Free Paid
Readability Layout1
Formula2
Bilingual3 Partial Partial
OCR4 7
Efficiency Batch tasks
Speed (sec/page) 1.47 1.50 1.86 1.67 0.38 1.88
  • 2
    Whether the tool can preserve figures, images, and formulas (although the formulas may be incorrectly displayed).
  • 2
    Whether the tool can translate documents containing formulas and correctly handle the positions of the formulas.
  • 3
    Whether the tool supports bilingual exports where both the original and translated text are simultaneously readable.
  • 4
    Whether the tool can handle scanned documents instead of digital documents.
  • 5
    IMT: Immersive Translate PDF Pro.
  • 6
    The original LaTeX file is required for translation.
  • 7
    Available in a community fork: PDFMathTranslate/PDFMathTranslate-next.
Table 1: Comparison with other projects and productions
Our comparison indicates that PDFMathTranslate surpasses alternative solutions. First, regarding accessibility, we significantly reduce user costs by providing an open-source, self-deployable, and free solution with flexible API integration. This flexibility enables broad applicability of our tool and ensures accurate translation outputs from any preferred service or model.Second, in terms of readability, our approach preserves basic layouts, accurately recognizes complex mathematical formulas, and supports bilingual outputs. Finally, although our solution is not the fastest—due to the computational demands of on-device inference—its efficiency remains comparable to other services (see 1).These advantages suggest that our tool becomes a better alternative to existing tools in the translation of scientific documents, thus potentially better reducing the language barriers.
4 Use Cases
We illustrate two typical use cases of PDFMathTranslate to illustrate how the tool translates text while preserving information embedded in layouts. The first use case is a technical document with both text and complex formulas, and the second use case is a research article with complex layouts and figures (see Figure 4).
Both cases show that the tool is capable of precisely translating the information within text while preserving the crucial information embedded in layouts.
Figure 4: Use cases.
5 Comparison with Existing Systems
To illustrate the advantages of our tool over other commercial and open-source alternatives in several important aspects, including accessibility, readability, and efficiency.
6 Sustainability
Nevertheless, maintenance is challenging in open-source projects Stol and Ali Babar . To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
First, we increase developer exposure by prominently featuring contributors on the project’s homepage via third-party services and regularly highlighting recent contributions in project documentation (e.g., 1970-01-01 / Supports Google translation (by @author_handle). This strategy has successfully incentivized 44 global developers who collectively contributed tens of thousands of lines of code.
Second, through collaboration with a commercial partner, we provide sponsored rewards such as membership exchanges for active contributors. These 11 subscriptions sent out to developers have effectively encouraged consistent contributions, resulting in the resolution of over 485 user-reported issues (by April 2025). Such cooperation has enabled the integration of advanced academic and community-driven technologies, including ONNX model support Wybxc (2025) and Qwen ws051682 (2025), thus advancing cross-linguistic scientific communication.
This sustainable collaboration model propelled the project to the top of GitHub’s global trends for over a week, garnering more than 25k stars, 22k downloads, and over 49k Docker pulls (by June 2025). Additionally, our project’s success has inspired a range of new commercial products with similar functionalities, significantly impacting both the scientific community and the broader public.
7 Limitations
While we have made several key improvements, our tool still has specific limitations. First, translation quality remains highly dependent on the underlying translation models and prompts. Second, the accuracy of the layout detection model in identifying various document layouts similarly depends on the quality of the layout detection models employed. Third, the current version of the tool can not yet handle scanned PDFs lacking optimal optical character recognition (OCR); however, this last feature is planned for implementation.
8 Acknowledgments
The authors thank all 44 contributors [^5] They are 7shi, Byaidu, Copilot, Cycloctane, Hanaasagi, IuvenisSapiens, JEFF-dev-ui, Tql-ws1, Wybxc, YadominJinta, Zxis233, alohays, aseaday, awwaawwa, borcation, charles7668, chiu0602, czz404, damaoooo, dependabot[bot], domonnss, eltociear, hellofinch, highkay, hotwa, imClumsyPanda, kharkover, kidach1, lintian233, mydreamworldpolly, namazuchin, qqueing, reycn, tastelikefeet, timelic, treeleaves30760, tylzh97, ws051682, wx-11, xxnuo, xyzyx233, yidasanqian, ymattw, zqqian; ordered alphabetically, retrieved through the official GitHub API by March 2025) of the open-sourced project on GitHub; platforms allowing us to host demos (Hugging Face, ModelScope); services we used in the demo (Google Translate); and all comments from users.
9 Ethics and Broader Impacts
Our work has already made a significant positive impact within both multilingual scientific and open-source communities. However, as a flexible, widely used tool, it raises potential ethical concerns regarding the copyright of documents. Translating documents without proper permissions could challenge the intellectual property rights of scientific works and innovations. We welcome suggestions from experts in intellectual property to mitigate the concern.
Appendix A Appendix
A.1 Appendix 1. Language and services supported in PDFMathTranslate
Table 2: Supported languages and services for translation in PDFMathTranslate
Category Details Total
Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 a
models Ollama Xinference, Gemma, OpenAI, OpenAI-likea, DeepSeek,
AzureOpenAI, Zhipu, ModelScope, Silicon, Gemini, Azure,
Tencent, Dify, AnythingLLM, Argos Translate, Grok, Groq
Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 b
output Portuguese (Brazil), Bulgarian, Catalan, Cherokee, Croatian,
languages Czech, Danish, Dutch, English (US), Estonian, Filipino,
Finnish, French, German, Greek, Gujarati, Hebrew, Hindi,
Hungarian, Icelandic, Indonesian, Italian, Japanese,
Kannada, Korean, Latvian, Lithuanian, Malay, Malayalam, Marathi,
Norwegian, Polish, Portuguese (Portugal), Romanian, Russian,
Serbian, Chinese (Simplified), Slovak, Slovenian, Spanish, Swahili,
Swedish, Tamil, Telugu, Thai, Chinese (Traditional), Turkish,
Urdu, Ukrainian, Vietnamese, Welsh
  • a
    Note. a The number of supported languages far exceeds what is shown in the table. Our protocol, designed as a general framework for OpenAI-like services, allows users to integrate any translation system, including locally hosted models or private online services, ensuring extensive language support.
  • b
    b Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.

### v2

@@
 Leveraging the most recent advances in large language models and precise layout detection, we contribute to the community with key improvements in precision, flexibility, and efficiency.
-The work has been open-sourced at https://github.com/byaidu/pdfmathtranslate with more than 22k downloads.
+The work is open-sourced at https://github.com/byaidu/pdfmathtranslate with more than 222k downloads.
 PDFMathTranslate: Scientific Document Translation Preserving Layouts
@@
 However, machine translation literature fails to address the unique challenges posed by the layouts in technical translation Schubert (2012).
-In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures have rich and important meanings.
+In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures has rich and important meanings.
 Those text-based attempts ignore the information in the layout and are thus insufficient to address the barrier in scientific and technical document translation.
@@
 To fill the gap, we introduce PDFMathTranslate, the world’s first open-source tool to translate PDF documents with preserved layouts.
-By leveraging recent advances in layout detection and large language models (see Figure 1), we better address the barrier with at least five key contributions: (1) efficient workflow of layout detection, translation, and re-rendering; (2) support for multiple languages (3) support for multiple translation models and services; (4) diverse user interfaces; and (5) a community-commerce model affording sustainable developments.
+By leveraging recent advances in layout detection and large language models (see Figure 1), we better address the barrier with at least five key contributions: (1) efficient workflow of layout detection, translation, and re-rendering; (2) support for multiple languages; (3) support for multiple translation models and services; (4) diverse user interfaces; and (5) a community-commerce model affording sustainable developments.
 2 Architecture and Design
-Overall, PDFMathTranslate is designed to translate documents while preserving their original layouts.
+PDFMathTranslate is designed to translate documents while preserving their original layouts.
 First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service (see Appendix A).
@@
 Technically, we design the architecture for precision, flexibility, and efficiency.
-Specifically, we introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, to be explained in the following three sub-sections.
+We introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, which we will explain in the following three subsections.
 Precise layout parser
-Preserving layouts prerequires precise parsing.
-To precisely parse the layouts, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering.
+Preserving layouts requires precise layout extraction.
+To precisely parse the position of document elements, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering.
 In the beginning, we exploit a recent advance in layout detection,  DocLayout-YOLO-DocStructBench-onnx.
@@
 To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX) onnx standard and a pytorch version.
-The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardwares.
+The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardware.
 Flexible translation middleware
@@
 Efficient streaming flow
-Translating scientific and technical documents, in some cases, needs an efficient workflow for thousands of pages, challenging the performance.
+Translating scientific and technical documents often necessitates an efficient workflow, especially when dealing with thousands of pages, which can challenge overall performance.
 To address the issue, we landed on a streaming design that is centered around an efficient in-memory processing pipeline that minimizes disk I/O while preserving document fidelity.
@@
 Developers who prefer command-line tools can conveniently install our Python CLI program using pip install pdf2zh.
-The command-line tool supports extensive features, with the major ones shown in Figure 3 and documented in detail  [^4] See https://github.com/Byaidu/PDFMathTranslate/blob
+The command-line tool supports extensive features, with the major options shown in Figure 3 and more documented in details  [^4] See https://github.com/Byaidu/PDFMathTranslate/blob
 /main/docs/ADVANCED.md
 .
-For users aiming to deploy applications on servers, we offer a Docker image designed specifically to simplify online production deployment.
-Those images are built and distributed on different platforms for developers with potential network issues.
+For users who want to deploy the software on servers, we offer a Docker image for simplified deployment.
+Additionally, Docker images are built and distributed on different platforms for developers with potential network issues.
 docker pull byaidu/pdf2zh
@@
 Speed (sec/page) 1.47 1.50 1.86 1.67 0.38 1.88
-  • 2
+  • 1
     Whether the tool can preserve figures, images, and formulas (although the formulas may be incorrectly displayed).
@@
 Table 1: Comparison with other projects and productions
-Our comparison indicates that PDFMathTranslate surpasses alternative solutions.
-First, regarding accessibility, we significantly reduce user costs by providing an open-source, self-deployable, and free solution with flexible API integration.
-This flexibility enables broad applicability of our tool and ensures accurate translation outputs from any preferred service or model.Second, in terms of readability, our approach preserves basic layouts, accurately recognizes complex mathematical formulas, and supports bilingual outputs.
-Finally, although our solution is not the fastest—due to the computational demands of on-device inference—its efficiency remains comparable to other services (see 1).These advantages suggest that our tool becomes a better alternative to existing tools in the translation of scientific documents, thus potentially better reducing the language barriers.
+4 Comparison with Existing Systems
+Compared with existing alternatives, PDFMathTranslate offers a superior solution from three perspectives: accessibility, readability, and efficiency (see Table 1).
+In terms of accessibility, the tool is open-sourced, self-deployable, supports API calls, and is entirely free of charge.
+These advantages ensure wider accessibility compared to any other existing solutions.
+Regarding readability, the tool preserves layouts, supports complex formulas, and is capable of exporting bilingual documents.
+Notably, a current limitation of our tool is its lack of optimal optical character recognition (OCR) support, which restricts its wider application for scanned documents.
+However, OCR functionality is planned for implementation in upcoming versions.
+Finally, concerning efficiency, while this application is slower than text-based translation services (like Google Translate) due to the integration of precise layout detection models, our solution is significantly faster than alternative products that also preserve layouts (such as IMT and Doc2X).
 5 Use Cases
 We illustrate two typical use cases of PDFMathTranslate to illustrate how the tool translates text while preserving information embedded in layouts.
-The first use case is a technical document with both text and complex formulas, and the second use case is a research article with complex layouts and figures (see Figure 4).
-Both cases show that the tool is capable of precisely translating the information within text while preserving the crucial information embedded in layouts.
+The first use case (Panel A, Figure 4) is an excerpt from a textbook with both text and complex formulas, and the second case (Panel B, Figure 4) is a scientific research article with complex layouts and figures (see Figure 4).
+Both cases consistently show that our work is capable of precisely translating the information within text while preserving the crucial information embedded in layouts.
 Figure 4: Use cases.
-5 Comparison with Existing Systems
-To illustrate the advantages of our tool over other commercial and open-source alternatives in several important aspects, including accessibility, readability, and efficiency.
 6 Sustainability
-Nevertheless, maintenance is challenging in open-source projects Stol and Ali Babar .
+Nevertheless, maintenance is challenging to open-source projects Stol and Ali Babar .
 To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
@@
 This strategy has successfully incentivized 44 global developers who collectively contributed tens of thousands of lines of code.
-Second, through collaboration with a commercial partner, we provide sponsored rewards such as membership exchanges for active contributors.
+Second, we provide sponsored rewards such as membership exchanges for active contributors through collaboration with a commercial partner.
 These 11 subscriptions sent out to developers have effectively encouraged consistent contributions, resulting in the resolution of over 485 user-reported issues (by April 2025).
 Such cooperation has enabled the integration of advanced academic and community-driven technologies, including ONNX model support Wybxc (2025) and Qwen ws051682 (2025), thus advancing cross-linguistic scientific communication.
-This sustainable collaboration model propelled the project to the top of GitHub’s global trends for over a week, garnering more than 25k stars, 22k downloads, and over 49k Docker pulls (by June 2025).
+This sustainable collaboration model propelled the project to the top of GitHub’s global trends for over a week, garnering more than 25k stars, 222k downloads, and over 49k Docker pulls (by June 2025).
 Additionally, our project’s success has inspired a range of new commercial products with similar functionalities, significantly impacting both the scientific community and the broader public.
@@
 Second, the accuracy of the layout detection model in identifying various document layouts similarly depends on the quality of the layout detection models employed.
-Third, the current version of the tool can not yet handle scanned PDFs lacking optimal optical character recognition (OCR); however, this last feature is planned for implementation.
+Third, the current version of the tool can not handle scanned PDFs lacking optimal optical character recognition (OCR); however, this last feature is planned for implementation.
 8 Acknowledgments
@@
 Appendix A Appendix
-A.1 Appendix 1. Language and services supported in PDFMathTranslate
+A.1 Officially supported language and services
 Table 2: Supported languages and services for translation in PDFMathTranslate
 Category Details Total
-Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 a
-models Ollama Xinference, Gemma, OpenAI, OpenAI-likea, DeepSeek,
+Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 1
+models Ollama Xinference, Gemma, OpenAI, OpenAI-like 1, DeepSeek,
 AzureOpenAI, Zhipu, ModelScope, Silicon, Gemini, Azure,
 Tencent, Dify, AnythingLLM, Argos Translate, Grok, Groq
-Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 b
+Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 2
 output Portuguese (Brazil), Bulgarian, Catalan, Cherokee, Croatian,
@@
 Urdu, Ukrainian, Vietnamese, Welsh
-  • a
-    Note.
-a The number of supported languages far exceeds what is shown in the table.
+  • 1
+    1 The number of supported languages far exceeds what is shown in the table.
 Our protocol, designed as a general framework for OpenAI-like services, allows users to integrate any translation system, including locally hosted models or private online services, ensuring extensive language support.
-  • b
-    b Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.
+  • 2
+    2 Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.

### v4

@@
 1 Introduction
-The advances in science and technologies spread and evolve in different languages  Von Gizycki ; Montgomery .
+Science and technologies diffuse and develop in different languages  Von Gizycki (1973); Montgomery (2013).
 Yet, language barriers hinder the diffusion and development of scientific progress Ramírez-Castañeda (2020); Hwang (2005); Ammon (2012).
-For example, while 98% publications in science are written in English Liu ; Ammon , the language only has 7.3% native speakers and no more than 20% speakers worldwide Bahji et al.
-.
-Consequently, the majority of human beings are hindered by language barriers from the advances in science and technology.
-To overcome the barrier, there are attempts such as those from international organizations UNESCO and academic efforts in improving Machine Translation (MT) over text Brown et al.
-(a); Vaswani et al.
+For example, while 98% publications in science are written in English Liu ; Ammon (2012), the language only has 7.3% native speakers and no more than 20% speakers worldwide Bahji et al.
+(2023).
+Consequently, the majority of human beings are hindered by language barriers from the advances in science and technologies.
+To overcome the barrier, there are the attempts from international organizations for inclusive science  [^1] See https://www.unesco.org/en/open-science and the attempts from academia in improving Machine Translation (MT) over text Brown et al.
+(1990); Vaswani et al.
 (2017); Zhu et al.
-; Johnson et al.
-; Sennrich et al.
-.
-However, machine translation literature fails to address the unique challenges posed by the layouts in technical translation Schubert (2012).
-In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures has rich and important meanings.
-Those text-based attempts ignore the information in the layout and are thus insufficient to address the barrier in scientific and technical document translation.
+(2020); Johnson et al.
+(2017); Sennrich et al.
+(2016).
+However, text-based machine translation fails to address the unique challenges posed by the layouts in technical translation Schubert (2012) where the elaborate access structure matters (pp.
+351-352).
+Non-textual elements in scientific and technical documents are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures have rich and important meanings.
+Thus, text-based machine translation ignoring such important information appear insufficient to address the barrier in scientific and technical document translation, hindering the progress of science, technologies, and society.
 Figure 1: The architecture of PDFMathTranslate
@@
 PDFMathTranslate is designed to translate documents while preserving their original layouts.
-First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service (see Appendix A).
+First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service.
 Next, it detects the layouts to extract the layouts and textual contents from the document.
@@
 Technically, we design the architecture for precision, flexibility, and efficiency.
-We introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, which we will explain in the following three subsections.
-Precise layout parser
+These three principles are realized by a precise parser, a flexible translation middle-ware, and an efficient workflow, detailed in following subsections.
+2.1 Precise layout parser
 Preserving layouts requires precise layout extraction.
@@
 The model is another version of a SOTA solution in object detection, Yolov10 Wang et al.
-, and is fast and accurate in the specific task Zhao et al.
-.
-To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX) onnx standard and a pytorch version.
+(2024), and is fast and accurate in the specific task Zhao et al.
+(2024).
+To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX)  [^2] See https://onnx.ai/ standard and a pytorch version.
 The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardware.
-Flexible translation middleware
+2.2 Flexible translation middleware
 Translation middleware offers flexibility in terms of language diversity, supported services, and customization capabilities.
-One, PDFMathTranslate supports at least 56 languages (see Appendix A), including popular ones such as English, and some languages shared by relatively smaller communities, such as Cherokee.
+One, PDFMathTranslate supports at least 194 languages (see Appendix A), including popular ones such as English, and some languages shared by relatively smaller communities, such as Cherokee.
 Users can freely select any of these languages as either the input or output language when translating documents.
@@
 This feature allows users to exploit advanced prompting strategies such as few-shots Brown et al.
-(b), Chain-of-Thought Wei et al.
-, Role-playing Shanahan et al.
-, etc.
-Efficient streaming flow
+(2020), Chain-of-Thought Wei et al.
+(2022), Role-playing Kong et al.
+(2024); Shanahan et al.
+(2023), etc.
+2.3 Efficient streaming flow
 Translating scientific and technical documents often necessitates an efficient workflow, especially when dealing with thousands of pages, which can challenge overall performance.
@@
 6 Sustainability
-Nevertheless, maintenance is challenging to open-source projects Stol and Ali Babar .
+Nevertheless, maintenance is challenging to open-source projects Stol and Ali Babar (2010).
 To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
-First, we increase developer exposure by prominently featuring contributors on the project’s homepage via third-party services and regularly highlighting recent contributions in project documentation (e.g., 1970-01-01 / Supports Google translation (by @author_handle).
+First, we increase developer exposure by prominently featuring contributors on the project’s homepage and regularly highlighting recent contributions (e.g., 1970-01-01 / Supports Google translation (by @author_handle).
 This strategy has successfully incentivized 44 global developers who collectively contributed tens of thousands of lines of code.
@@
 8 Acknowledgments
-The authors thank all 44 contributors [^5] They are 7shi, Byaidu, Copilot, Cycloctane, Hanaasagi, IuvenisSapiens, JEFF-dev-ui, Tql-ws1, Wybxc, YadominJinta, Zxis233, alohays, aseaday, awwaawwa, borcation, charles7668, chiu0602, czz404, damaoooo, dependabot[bot], domonnss, eltociear, hellofinch, highkay, hotwa, imClumsyPanda, kharkover, kidach1, lintian233, mydreamworldpolly, namazuchin, qqueing, reycn, tastelikefeet, timelic, treeleaves30760, tylzh97, ws051682, wx-11, xxnuo, xyzyx233, yidasanqian, ymattw, zqqian; ordered alphabetically, retrieved through the official GitHub API by March 2025) of the open-sourced project on GitHub; platforms allowing us to host demos (Hugging Face, ModelScope); services we used in the demo (Google Translate); and all comments from users.
+The authors thank all 44 contributors [^7] They are 7shi, Byaidu, Copilot, Cycloctane, Hanaasagi, IuvenisSapiens, JEFF-dev-ui, Tql-ws1, Wybxc, YadominJinta, Zxis233, alohays, aseaday, awwaawwa, borcation, charles7668, chiu0602, czz404, damaoooo, dependabot[bot], domonnss, eltociear, hellofinch, highkay, hotwa, imClumsyPanda, kharkover, kidach1, lintian233, mydreamworldpolly, namazuchin, qqueing, reycn, tastelikefeet, timelic, treeleaves30760, tylzh97, ws051682, wx-11, xxnuo, xyzyx233, yidasanqian, ymattw, zqqian; ordered alphabetically, retrieved through the official GitHub API by March 2025) of the open-sourced project on GitHub; platforms allowing us to host demos (Hugging Face, ModelScope); services we used in the demo (e.g., Google Translate); redeem codes from Immersive Translate; public API access from SiliconFlow; and all comments from users.
 9 Ethics and Broader Impacts
-Our work has already made a significant positive impact within both multilingual scientific and open-source communities.
+Our work has made a significant positive impact within both multilingual scientific and open-source communities.
 However, as a flexible, widely used tool, it raises potential ethical concerns regarding the copyright of documents.
@@
 A.1 Officially supported language and services
+In Table 2, we enumerate the officially supported services, models, and languages. However, the actual number may be significantly higher due to the extensibility of services and models.
 Table 2: Supported languages and services for translation in PDFMathTranslate
 Category Details Total
-Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 1
-models Ollama Xinference, Gemma, OpenAI, OpenAI-like 1, DeepSeek,
-AzureOpenAI, Zhipu, ModelScope, Silicon, Gemini, Azure,
+Service or Google (default), 302.AI, Bing, DeepL, DeepLX, Ollama, Ali Qwen, 23 1
+models Ollama, Xinference, Gemma, OpenAI, OpenAI-like 1, DeepSeek,
+AzureOpenAI, Zhipu, ModelScope, Silicon Cloud, Gemini, Azure,
 Tencent, Dify, AnythingLLM, Argos Translate, Grok, Groq
-Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 2
-output Portuguese (Brazil), Bulgarian, Catalan, Cherokee, Croatian,
-languages Czech, Danish, Dutch, English (US), Estonian, Filipino,
-Finnish, French, German, Greek, Gujarati, Hebrew, Hindi,
-Hungarian, Icelandic, Indonesian, Italian, Japanese,
-Kannada, Korean, Latvian, Lithuanian, Malay, Malayalam, Marathi,
-Norwegian, Polish, Portuguese (Portugal), Romanian, Russian,
-Serbian, Chinese (Simplified), Slovak, Slovenian, Spanish, Swahili,
-Swedish, Tamil, Telugu, Thai, Chinese (Traditional), Turkish,
-Urdu, Ukrainian, Vietnamese, Welsh
+Input and Abkhaz, Acehnese, Acholi, Afrikaans, Albanian, Alur, Amharic, 194 2
+output Arabic, Armenian, Assamese, Awadhi, Aymara, Azerbaijani, Balinese,
+languages Bambara, Bashkir, Basque, Batak Karo, Batak Simalungun,
+Batak Toba, Belarusian, Bemba, Bengali, Betawi, Bhojpuri,
+Bikol, Bosnian, Breton, Bulgarian, Buryat, Cantonese,
+Catalan, Cebuano, Chichewa (Nyanja), Chinese (Simplified),
+Chinese (Traditional), Chuvash, Corsican, Crimean Tatar,
+Croatian, Czech, Danish, Dinka, Divehi, Dogri, Dombe,
+Dutch, Dzongkha, English, Esperanto, Estonian, Ewe,
+Fijian, Filipino (Tagalog), Finnish, French,
+French (French), French (Canadian), Frisian,
+Fulfulde, Ga, Galician, Ganda (Luganda), Georgian,
+German, Greek, Guarani, Gujarati, Haitian Creole,
+Hakha Chin, Hausa, Hawaiian, Hebrew, Hiligaynon,
+Hindi, Hmong, Hungarian, Hunsrik, Icelandic, Igbo,
+Iloko, Indonesian, Irish, Italian, Japanese, Javanese,
+Kannada, Kapampangan, Kazakh, Khmer, Kiga, Kinyarwanda,
+Kituba, Konkani, Korean, Krio, Kurdish (Kurmanji),
+Kurdish (Sorani), Kyrgyz, Lao, Latgalian, Latin, etc.
   • 1
@@
   • 2
-    2 Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.
+    2 Based on the default service (Google Translate) and interface (CLI).